# Yogesh Kulkarni, Pooyan Fazli

### v1

@@ Abstract
Multimodal reasoning over long-horizon video is challenging due to the need for precise spatiotemporal fusion and alignment across modalities. While recent methods such as Group Relative Policy Optimization (GRPO) have shown promise in this domain, they suffer from three key limitations: (1) data inefficiency from their on-policy design, (2) a vanishing advantage problem, where identical or near-identical rewards within a group eliminate the learning signal by producing zero-valued advantages, and (3) uniform credit assignment that fails to emphasize critical reasoning steps. We introduce AVATAR (Audio-Video Agent for Alignment and Reasoning), a framework that addresses these limitations through two core components: (1) an off-policy training architecture that improves sample efficiency and resolves vanishing advantages by reusing past experiences with greater reward diversity, and (2) Temporal Advantage Shaping (TAS), a novel credit assignment strategy that upweights key reasoning phases during learning. AVATAR achieves strong performance across various benchmarks, outperforming the Qwen2.5-Omni baseline by +5.4 on MMVU, +4.9 on OmniBench, and +4.5 on Video-Holmes, while demonstrating over 35% higher sample efficiency.
Introduction
@@
Multimodal large language models (MLLMs) must align video, audio, and language modalities to support long-horizon reasoning (Lu et al. 2025). This requires balancing dense temporal coverage for narrative understanding with high spatial resolution for visual grounding (Zhong et al. 2025; Chen et al. 2025b). Reinforcement learning (RL), particularly Group Relative Policy Optimization (GRPO)(DeepSeek-AI et al. 2025; Chen et al. 2025a), has emerged as a promising approach for enhancing such reasoning. However, applying GRPO directly to multimodal domains reveals key limitations.
@@
Figure 1: Standard GRPO (top) vs. AVATAR (bottom). AVATAR enhances standard GRPO with two key components: (1) an off-policy architecture using a stratified replay buffer to improve data efficiency, and (2) Temporal Advantage Shaping (TAS), a novel credit assignment strategy that focuses learning on critical reasoning steps.
@@
First, GRPO is an on-policy method, discarding samples after each update. This makes it data-inefficient, a critical issue in the video domain where learning models rely on expensive, large-scale annotated data (Liu et al. 2025c; Wang et al. 2025b, e). Prior methods like RePO (Li et al. 2025c) reuse past data via replay buffers but lack selective control over replay quality. Second, GRPO suffers from the vanishing advantage problem: when the reward variance within a group collapses, such as in cases where all responses are either uniformly correct or uniformly incorrect, the resulting zero-valued advantage signals prevent effective learning (Wang et al. 2025a; Liu et al. 2025a; Dai, Yang, and Si 2025). Third, GRPO suffers from uniform credit assignment (Samineni et al. 2025; Deng et al. 2025), assigning the same reward to all tokens in a reasoning chain, regardless of their contribution to the final response (Dang et al. 2025; Fan et al. 2025a; Meng et al. 2025). This issue is amplified in multimodal tasks, where the GRPO objective fails to distinguish visually grounded reasoning from plausible text. By treating fluent language and accurate perception as equivalent, it enables the model to produce correct-sounding responses that ignore visual input, resulting in frequent perception errors (Wang et al. 2025d; Fan et al. 2025b).
@@
To address these challenges, we introduce AVATAR (Audio-Video Agent for Alignment and Reasoning), a unified framework built on GRPO to enhance multimodal reasoning, with a focus on video tasks. To address GRPO’s data inefficiency, AVATAR adopts an off-policy training architecture that reuses past experiences through a stratified replay buffer. Samples are grouped based on how often the policy has succeeded on them in the past. This creates a curriculum that helps the model focus on samples it previously failed on, which is especially important for learning from sparse and challenging audio-video data. Moreover, to improve training stability and reward signal quality, we extend prior clue-grounded rewards (Lu et al. 2025; Lai et al. 2025; Huang et al. 2025) that check whether the model uses relevant audio or visual cues from the query. We introduce two new components: stepwise rewards, which give partial credit for intermediate reasoning, and a Video-Context Reference Score (VCRS), which uses a moving average of recent rewards for each query to maintain a stable and non-zero advantage signal. We also address the uniform credit assignment problem with a new method called Temporal Advantage Shaping (TAS). Unlike prior work that treats all tokens equally (Samineni et al. 2025) or depends on statistical variance (Dang et al. 2025), TAS guides learning based on a token’s position in the reasoning chain, placing greater weight on the beginning (planning) and end (synthesis) stages. This is especially useful for audio-video tasks, where accurate early grounding (e.g., locating the speaker) and final synthesis (e.g., combining speech and visual cues to identify the speaker) are critical. In the experiments, we apply AVATAR within a three-stage RL training pipeline to evaluate its effectiveness: starting with general visual reasoning, followed by audio-visual alignment, and concluding with fine-grained audio-based object localization. In summary, our primary contributions are:
@@
  • We present AVATAR, an off-policy RL framework that enhances data efficiency in audio-video reasoning by leveraging past experiences through a difficulty-aware replay buffer, which focuses on challenging samples.
@@
  • We introduce new approaches for RL reward signaling and credit assignment by giving partial credit to intermediate reasoning steps, stabilizing learning with a moving average of past rewards, and emphasizing the most critical parts of the reasoning process.
@@
  • Experiments show that AVATAR achieves significant gains over the Qwen2.5 Omni baseline, with improvements of +5.4 on MMVU, +4.9 on OmniBench, and +4.5 on Video-Holmes, while demonstrating over 35% higher sample efficiency.
Related Work
@@
Recent MLLMs have made remarkable progress towards audio-video understanding (Bai et al. 2025; Zhu et al. 2025; Sun et al. 2025; Fu et al. 2024; Tang et al. 2025; Islam et al. 2025). Key efforts focus on architectural enhancements like dedicated audio branches and video connectors (Cheng et al. 2024), using clue aggregators to ground responses in question-specific details (Ye et al. 2024), or with token interleaving to enhance temporal understanding (Tang et al. 2025). However, these methods excel at direct video question answering but cannot rationalize their thoughts through reasoning traces. This limitation is problematic for complex audio-visual tasks requiring multi-step reasoning and fine-grained cross-modal alignment, such as localizing sound sources. In contrast, AVATAR enhances structured reasoning capabilities in audio-visual models using a sample-efficient RL approach.
@@
Early multimodal efforts used preference optimization like DPO (Rafailov et al. 2023; Kulkarni and Fazli 2025a, b; Sarkar and Etemad 2025), but RL methods such as GRPO (DeepSeek-AI et al. 2025; Chen et al. 2025a) have shown greater promise for reasoning. Recent variants adapt GRPO to specific tasks: Video-R1 uses temporal contrastive rewards (Feng et al. 2025), HumanOmniV2 employs LLM-judged rewards for global audio-visual context (Zhao et al. 2025a), Omni-R1 targets audio-visual segmentation (Zhong et al. 2025), and GRIT generates spatial reasoning chains (Fan et al. 2025b). However, these approaches suffer from uniform credit assignment and sample inefficiency. AVATAR addresses these limitations through its off-policy architecture and TAS credit assignment. For audio-visual tasks, it’s U-shaped temporal weighting emphasizes early grounding (localizing sounds) and late synthesis (combining auditory and visual evidence), while the stratified replay buffer handles sparse cross-modal data effectively.
Preliminaries
Group Relative Policy Optimization
GRPO is an RL framework that can be used to fine-tune MLLMs by comparing multiple candidate responses generated by a behavior policy $\pi_{\theta_{\text{old}}}$ for a given prompt $q$ (DeepSeek-AI et al. 2025). For each group of $K$ candidate responses $\{o_{i}\}_{i=1}^{K}$, scalar rewards $R(o_{i})$ from predefined reward functions are assigned to each response. These rewards are then normalized to compute relative advantages $A_{i}$ within the group.
$A_{i}=\frac{R(o_{i})-\mu_{R}}{\sigma_{R}+\epsilon_{\text{adv}}},$ (1)
where $\mu_{R}$ and $\sigma_{R}$ are the mean and standard deviation of the group rewards $\{R(o_{i})\}_{i=1}^{K}$ and $\epsilon_{adv}$ is a small constant for numerical stability. These group-relative advantages guide the update of the target policy $\pi_{\theta}$, encouraging it to assign higher probability to relatively better responses. GRPO maximizes the following clipped surrogate objective with KL regularization (Li et al. 2025c):
$\begin{split}\mathcal{J}_{\text{GRPO}}(\theta)&=\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}\sim\pi_{\theta_{old}}}\left[\frac{1}{K}\sum_{i=1}^{K}\min(r_{i}A_{i},\right.\\ &\quad\left.\text{clip}(r_{i},1-\epsilon,1+\epsilon)A_{i})-\beta D_{KL}(\pi_{\theta}\|\pi_{\text{ref}})\right],\end{split}$ (2)
where $q$ is a prompt sampled from the dataset $\mathcal{D}$, $r_{i}$ is the importance sampling ratio $\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}$, $\epsilon$ controls the clipping range, and $\beta$ weights the KL divergence term to ensure stability by regularizing the updated policy $\pi_{\theta}$ against a reference policy $\pi_{\text{ref}}$.
@@
GRPO is an on-policy method that updates using only fresh rollouts from the current policy and discards past data. This limits sample reuse and leads to inefficiency, especially in video understanding tasks where data is costly to collect. Off-policy updates offer a more efficient alternative by reusing past experiences to reduce training costs.
@@
A fundamental limitation of GRPO arises from its advantage estimation mechanism. When the rewards $R(o_{i})$ for all outputs in a group are identical or nearly identical (e.g., all zero for a particularly difficult prompt or all one for an easy prompt), the numerator of the advantage function, $R(o_{i})-\mu_{R}$, becomes zero for all samples $i$. This results in zero advantages $(A_{i}=0)$, eliminating the learning signal and preventing the policy from updating. Consequently, GRPO fails to learn from such training samples, a phenomenon referred to as the vanishing advantage problem (Wang et al. 2025a; Li et al. 2025c).
@@
Another key limitation of GRPO is its overly simplistic credit assignment. In Equation 2, a single scalar advantage $A_{i}$ is computed for the entire output $o_{i}$ and applied uniformly to all tokens during policy updates. This ignores the varying importance of different reasoning steps. Tokens critical to initial planning receive the same learning signal as those from less important phases, failing to address the temporal credit assignment challenge that RL is designed to solve.
 [Uncaptioned image] AVATAR
We introduce AVATAR, an audio-video agent for multimodal alignment and reasoning, designed to address key limitations of GRPO. To tackle the vanishing advantage problem and resulting data inefficiency, AVATAR employs an off-policy training architecture using a stratified replay buffer. To overcome uniform credit assignment, it introduces Temporal Advantage Shaping (TAS), which modulates advantages (i.e., re-weighs the learning signal for each token) across the reasoning sequence. The following sections detail these components.
@@
1:Input: Initial policy $\pi_{\theta_{init}}$, Reward functions $R$, Hyperparameters $\alpha,\lambda_{\text{TAS}},\beta,\epsilon,K_{on},K_{off}$.
2:Initialize: Policy $\pi_{\theta}\leftarrow\pi_{\theta_{init}}$, Reference policy $\pi_{\text{ref}}\leftarrow\pi_{\theta_{init}}$, empty Difficulty-Aware Buffer $\mathcal{B}$.
@@
4:  Sample a batch of prompts $Q_{batch}$ from dataset $\mathcal{D}$.
5:  $\pi_{\theta_{old}}\leftarrow\pi_{\theta}$ {Set behavior policy}
6:  {Experience Generation}
7:  Generate $K_{on}$ on-policy experiences $E_{on}=\{(q,o,R(o))\}$ using $\pi_{\theta_{old}}$ and compute advantages $A_{on}$.
8:  Sample $K_{off}$ off-policy experiences $E_{off}=\{(q,o,\pi_{\theta_{off}})\}$ from $\mathcal{B}$ and compute advantages $A_{off}$.
9:  {Temporal Advantage Shaping (TAS)}
10:  for each sample $o_{i}$ with advantage $A_{i}$ in $E_{on}\cup E_{off}$ do
11:   Compute per-token weights $w_{i,t}$ using Equation 6.
12:   Create token-level shaped advantage: $A^{\text{TAS}}_{i,t}\leftarrow A_{i}\cdot w_{i,t}$.
13:  end for
14:  {Policy Update}
15:  Compute total loss $\mathcal{J}_{\text{total}}(\theta)$ by combining the on-policy and off-policy surrogate objectives using their respective shaped advantages ($A^{\text{TAS}}_{on},A^{\text{TAS}}_{off}$) and the final KL penalty term.
16:  Update policy: $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{J}_{\text{total}}(\theta)$.
17:  {Buffer Update}
18:  Update Buffer $\mathcal{B}$ with the new on-policy experiences.
19:end for
Off-Policy Architecture
@@
Our framework uses a stratified replay buffer $\mathcal{B}$ divided into three dynamic tiers: Easy (25%), Medium (35%), and Hard (40%). Experiences are assigned based on the policy’s moving average reward $\bar{R}(q)$ for each prompt $q$, forming a progressive learning curriculum. Fixed tier capacities ensure rapid turnover of easy samples while retaining harder ones for repeated training.
Buffer Update and Hinting.
The buffer uses two metrics: the moving average reward $\bar{R}(q)$ to assign difficulty tiers, and KL-divergence $D_{KL}(\pi_{\theta}||\pi_{\beta})$ between the target policy and behavior policy to monitor policy stability. If a prompt remains hard (low $\bar{R}(q)$) and the policy stops exploring (low KL), a hint mechanism is triggered. Pre-computed hints are generated by providing the full problem context (video, audio caption, query, and ground truth) to Qwen2.5-VL (Bai et al. 2025) (e.g., “first locate the object making the sound, then count”) and are added to guide learning and escape local optima. This keeps training in a “challenging but solvable” zone, applying curriculum learning principles while enabling targeted exploration (Huang et al. 2025).
Off-Policy Update with Selective Sample Replay.
At each training step, we form a batch using both newly generated on-policy samples and off-policy samples from the replay buffer. To handle policy drift in off-policy data, we apply importance sampling. The full objective function for AVATAR combines the standard on-policy GRPO loss with a corrected off-policy term:
$\mathcal{J}_{\text{AVATAR}}=\mathcal{J}_{\text{on-policy}}(\theta)+\alpha\cdot\mathcal{J}_{\text{off-policy}}(\theta),$ (3)
Where $\alpha$ is a hyperparameter balancing the off-policy contribution, and the off-policy objective $\mathcal{J}{\text{off-policy}}$ follows Equation 2, with the importance sampling ratio $r_{i}$ computed using the generating policy $\pi{\theta_{\text{off}}}$ stored in the replay buffer:
@@
This hybrid training strategy, inspired by recent work on off-policy PPO and GRPO (Li et al. 2025c; Mroueh et al. 2025), enables AVATAR to be significantly more data-efficient, while the buffer ensures that the replayed samples provide a stable and informative learning signal. The training loop is detailed in Algorithm 1.
Temporal Advantage Shaping (TAS)
Solving complex tasks typically requires more reasoning effort during the initial planning and final synthesis stages than during intermediate evidence gathering. To address GRPO’s uniform credit assignment issue, we introduce Temporal Advantage Shaping (TAS), a method that adjusts each token’s learning signal based on its position in the reasoning chain. This refines the single-advantage formulation used in standard GRPO (Equation 2). As shown in Figure 2, TAS applies a U-shaped weighting curve that amplifies the contribution of tokens in the planning and synthesis phases, unlike the flat, uniform credit assignment used in standard GRPO. This prevents overemphasizing low-value tokens and better reinforces critical reasoning steps.
Figure 2: To address GRPO’s uniform credit assignment (gray line), TAS applies a U-shaped weight to amplify advantages during crucial planning and synthesis stages.
U-Shaped Weight Calculation.

### v2

# Yogesh Kulkarni  Pooyan Fazli

@@ Abstract
[…]
1 Introduction
@@
Multimodal large language models (MLLMs) must align video, audio, and language modalities to support long-horizon reasoning [avreasoner]. This requires balancing dense temporal coverage for narrative understanding with high spatial resolution for visual grounding [omnir1, scalingrl, vidhalluc]. Reinforcement learning (RL), particularly Group Relative Policy Optimization (GRPO) [deepseek, sftorrl], has emerged as a promising approach for enhancing such reasoning. However, applying GRPO directly to multimodal domains reveals key limitations.
@@
Figure 1: Standard GRPO (top) vs. AVATAR (bottom). AVATAR enhances GRPO with two key components: (1) an off-policy architecture using a stratified replay buffer to improve data efficiency, and (2) Temporal Advantage Shaping (TAS), a novel credit assignment strategy that focuses learning on critical reasoning steps.
@@
First, GRPO is an on-policy method, discarding samples after each update. This makes it data-inefficient, a critical issue in the video domain where learning models rely on expensive, large-scale annotated data [Liu_2025_ICCV, videorft, videorts]. Prior methods like RePO [repo] or standard training frameworks like veRL [verl] reuse past data but lack selective control over replay quality. Second, GRPO suffers from the vanishing advantage problem: when the reward variance within a group collapses, such as in cases where all responses are either uniformly correct or uniformly incorrect, the resulting zero-valued advantage signals prevent effective learning [vlrethinker, understandingr1zero, sgrpo, feng2025don]. Third, GRPO suffers from uniform credit assignment [rlonlyinname, nthr], assigning the same reward to all tokens in a reasoning chain, regardless of their contribution to the final response [twgrpo, sophia, videocap]. This issue is amplified in multimodal tasks, where the GRPO objective fails to distinguish visually grounded reasoning from plausible text. By treating fluent language and accurate perception as equivalent, the model can produce correct-sounding responses that ignore visual input, resulting in frequent perception errors [papo, grit].
@@
To address these challenges, we introduce AVATAR (Audio-Video Agent for Alignment and Reasoning), a unified framework built on GRPO to enhance multimodal reasoning, with a focus on video tasks. To mitigate GRPO’s data inefficiency, AVATAR adopts an off-policy training architecture that reuses past experiences through a stratified replay buffer. Samples are grouped based on how often the policy has succeeded on them in the past. This creates a curriculum that helps the model focus on samples it previously failed on, which is especially important for learning from sparse and challenging audio-video data. Moreover, to enhance training stability and the quality of the reward signal, we build upon prior clue-grounded reward formulations [avreasoner, medr1, hintgrpo], which evaluate whether the model leverages relevant audio or visual cues from the query.
We also address the uniform credit assignment problem with a new method called Temporal Advantage Shaping (TAS). Unlike prior work that treats all tokens equally [rlonlyinname] or depends on statistical variance [twgrpo], TAS guides learning based on a token’s position in the reasoning chain, placing greater weight on the beginning (planning) and end (synthesis) stages. This is especially useful for audio-video tasks, where accurate early grounding (e.g., locating the speaker) and final synthesis (e.g., combining speech and visual cues to identify the speaker) are critical. In the experiments, we apply AVATAR within a three-stage RL training pipeline to evaluate its effectiveness: starting with general visual reasoning, followed by audio-visual alignment, and concluding with fine-grained audio-based object localization. In summary, our primary contributions are:
  1. 1.
@@
    We present AVATAR, an off-policy RL framework designed to improve data efficiency in audio-video reasoning by leveraging past experiences through a difficulty-aware replay buffer that prioritizes challenging samples.
  2. 2.
@@
    We introduce new approaches for reinforcement learning reward signaling and credit assignment by allocating partial credit to intermediate reasoning steps, stabilizing learning through a moving average of past rewards, and emphasizing the critical components of reasoning.
  3. 3.
@@
    Experiments show that AVATAR achieves significant gains over the Qwen2.5 Omni baseline, with improvements of +5.4+5.4 on MMVU, +4.9+4.9 on OmniBench, and +4.5+4.5 on Video-Holmes, while requiring $80\%$ fewer generated completions to reach target performance.
2 Related Work
@@
Recent MLLMs have made remarkable progress towards audio-video understanding [qwen25vl, internvl3, salmonn, vita, empowering, bimba]. Key efforts focus on architectural enhancements such as dedicated audio branches and video connectors [videollama2], using clue aggregators to ground responses in question-specific details [cat], or with token interleaving to enhance temporal understanding [empowering]. However, these methods excel at direct video question answering but cannot rationalize their thoughts through reasoning traces. This limitation is problematic for complex audio-visual tasks requiring multi-step reasoning and fine-grained cross-modal alignment, such as localizing sound sources. In contrast, AVATAR enhances structured reasoning capabilities in audio-visual models using a sample-efficient RL approach.
@@
Early multimodal efforts use preference optimization like DPO [dpo, videopasta, videosavi, rrpo], but RL methods such as GRPO [deepseek, sftorrl] have shown greater promise for reasoning. Recent variants adapt GRPO to specific tasks: Video-R1 uses temporal contrastive rewards [videor1], HumanOmniV2 employs LLM-judged rewards for global audio-visual context [humanomni], Omni-R1 targets audio-visual segmentation [omnir1], and GRIT generates spatial reasoning chains [grit]. However, these approaches suffer from uniform credit assignment and sample inefficiency. AVATAR addresses these limitations through its off-policy architecture and TAS credit assignment. For audio-visual tasks, its temporal weighting emphasizes early grounding (localizing sounds) and late synthesis (combining auditory and visual evidence), while the stratified replay buffer handles sparse cross-modal data effectively.
3 Preliminaries
3.1 Group Relative Policy Optimization
GRPO is an RL framework that can be used to fine-tune MLLMs by comparing multiple candidate responses generated by a behavior policy $\pi_{\theta_{\text{old}}}$ for a given prompt $q$ [deepseek]. Each candidate output $o_{i}$ typically includes both intermediate reasoning steps and a final answer. For each group of $K$ candidate responses $\{o_{i}\}_{i=1}^{K}$, scalar rewards $R(o_{i})$ from predefined reward functions are assigned to each response. These rewards are then normalized within the group to obtain relative advantages $A_{i}$.
$A_{i}=\frac{R(o_{i})-\mu_{R}}{\sigma_{R}+\epsilon_{\mathrm{adv}}},$ (1)
where $\mu_{R}$ and $\sigma_{R}$ are the mean and standard deviation of the group rewards $\{R(o_{i})\}_{i=1}^{K}$ and $\epsilon_{adv}$ is a small constant for numerical stability. These group-relative advantages guide the update of the target policy $\pi_{\theta}$, encouraging it to assign higher probability to relatively better responses. GRPO maximizes the following clipped surrogate objective with KL regularization [repo]:
$\begin{split}\mathcal{J}_{\mathrm{GRPO}}(\theta)&=\mathbb{E}_{\begin{subarray}{c}q\sim\mathcal{D}\\ \{o_{i}\}_{i=1}^{K}\sim\pi_{\theta_{\mathrm{old}}}\end{subarray}}\Bigg[\frac{1}{K}\sum_{i=1}^{K}\min\!\big(r_{i}A_{i},\;\bar{r}_{i}A_{i}\big)\\ &\qquad-\beta\,D_{\mathrm{KL}}\!\big(\pi_{\theta}(\cdot|q)\,\|\,\pi_{\text{ref}}(\cdot|q)\big)\Bigg],\end{split}$ (2)
where $\bar{r}_{i}=\mathrm{clip}(r_{i},1-\epsilon,1+\epsilon)$, $q$ is a prompt sampled from the dataset $\mathcal{D}$, $r_{i}$ is the importance sampling ratio $\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}$, $\epsilon$ controls the clipping range, and $\beta$ weights the KL divergence term to ensure stability by regularizing the updated policy $\pi_{\theta}$ against a reference policy $\pi_{\mathrm{ref}}$. Here, $\pi_{\mathrm{ref}}$ refers to the pretrained model before RL fine-tuning, used to prevent large deviations from the original policy.
3.2 GRPO Limitations
@@
GRPO is an on-policy method that discards experiences after a single update, a highly inefficient design for audio-visual reasoning, which relies on scarce, costly annotated data. When the pre-trained policy struggles with a complex prompt, this on-policy nature forces it to “move on,” discarding the informative failure experience instead of learning from it. Off-policy updates offer a more efficient alternative by creating a memory of these past experiences, enabling repeated learning on the most challenging samples.
@@
A fundamental limitation of GRPO arises from its advantage estimation mechanism. When the rewards $R(o_{i})$ for all outputs in a group are identical or nearly identical (e.g., all zero for a particularly difficult prompt or all one for an easy prompt), the group mean baseline $\mu_{R}$ equals each reward. As a result, the advantages become zero for all samples $i$: $A_{i}=R(o_{i})-\mu_{R}=0$. This eliminates the learning signal and prevents the policy from updating. Consequently, GRPO fails to learn from such training samples, a phenomenon referred to as the vanishing advantage problem [vlrethinker, repo].
@@
Another key limitation of GRPO is its overly simplistic credit assignment. In Equation 2, a single scalar advantage $A_{i}$ is computed for the entire output $o_{i}$ and applied uniformly to all tokens during policy updates. This ignores the varying importance of different reasoning steps or temporal phases. Tokens critical to initial planning receive the same learning signal as those from less important phases, failing to address the temporal credit assignment challenge that RL is designed to solve.
4 AVATAR  [Uncaptioned image]
We introduce AVATAR, an off-policy RL framework for multimodal alignment and reasoning, designed to address key limitations of GRPO. To tackle the vanishing advantage problem and resulting data inefficiency, AVATAR employs an off-policy training architecture using a stratified replay buffer. To overcome uniform credit assignment, it introduces Temporal Advantage Shaping (TAS), which modulates advantages across the reasoning sequence by upweighting tokens at the beginning and end of the sequence, ensuring that critical planning and final decision steps receive stronger learning signals.
@@ 4 AVATAR  [Uncaptioned image]
1:Input: Initial policy $\pi_{\theta_{init}}$, Rewards $R$, Hyperparams $\alpha,\lambda_{\text{TAS}},\beta,\epsilon,K_{on},K_{off}$.
2:Initialize: $\pi_{\theta}\leftarrow\pi_{\theta_{init}}$, $\pi_{\text{ref}}\leftarrow\pi_{\theta_{init}}$, empty Buffer $\mathcal{B}$, VCRS $\overline{R}(q)\leftarrow\{\}$.
@@
4:  Sample prompts $Q_{batch}$; $\pi_{\theta_{old}}\leftarrow\pi_{\theta}$
5:  Generate $K_{on}$ on-policy $E_{on}\!=\!\{(q,o,R(o))\}$ using $\pi_{\theta_{old}}$; Compute $A_{on}$ (Eq. 1).
6:  Retrieve VCRS $\overline{R}(q)$; Sample $K_{off}$ off-policy $E_{off}\!=\!\{(q,o,R(o),\pi_{\theta_{off}})\}$ from $\mathcal{B}$; Compute $A_{off}$ using $\overline{R}(q)$ (Eq. 11).
7:  Compute shaped advantages $A^{\text{TAS}}_{i,t}\leftarrow(A_{on}\cup A_{off})\cdot w_{i,t}$ using TAS weights $w_{i,t}$ (Eq. 6).
8:  Compute $\mathcal{J}_{\text{total}}(\theta)$ using $A^{\text{TAS}}_{on},A^{\text{TAS}}_{off}$ and importance ratios $r_{i}^{on},r_{i}^{off}$ (Eq. 3).
9:  Update policy: $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{J}_{\text{total}}(\theta)$.
10:  Update $\mathcal{B}$ with $E_{on}$ (with stratification/hinting ); Update $\overline{R}(q)$ using $E_{on}$.
11:end for
4.1 Off-Policy Architecture
@@
We employ a stratified replay buffer $\mathcal{B}$ to improve sample efficiency via a progressive curriculum. The buffer of size $10k$ is divided into three fixed-capacity tiers based on difficulty: Easy ($25\%$), Medium ($35\%$), and Hard ($40\%$). At each training step, new experiences $(q,o,\pi_{\theta_{\text{off}}})$, required for off-policy importance sampling, are stored. Tier assignment is determined not by individual rewards, but by the prompt’s moving average reward $\bar{R}(q)$, which serves as a dynamic difficulty metric. The tier thresholds are dynamic quantiles based on the $\bar{R}(q)$ score distribution, assigning the bottom $40\%$ of experiences to ‘Hard’, the next $35\%$ to ‘Medium’, and the top $25\%$ to ‘Easy’. As the agent’s performance on a prompt $q$ improves, its $\bar{R}(q)$ rises, and new experiences for that prompt are promoted to easier tiers. This curriculum design ensures difficult samples (low $\bar{R}(q)$) are retained in the high-capacity ‘Hard’ tier, forcing the model to learn from its failure modes through repeated exposure.
Hinting Mechanism.
AVATAR leverages the KL divergence $D_{KL}(\pi_{\theta},|,\pi_{\beta})$ between the target policy $\pi_{\theta}$ and behavior policy $\pi_{\beta}$ to monitor policy stability. When a prompt remains hard (low $\bar{R}(q)$) and the policy stops exploring (low KL), a pre-computed hint is triggered. Hints (e.g., “first locate the object making the sound, then count”) are generated by providing the full problem context (video, audio caption, query, and ground truth) to Qwen2.5-VL [qwen25vl] and are used as additional guidance to help the agent escape local optima. This ensures training remains in a challenging but solvable regime while promoting targeted exploration [hintgrpo]. Prompt provided in Supplementary, Figure 13.
Hybrid Training with Selective Replay.
At each training step, a batch is formed from both newly generated on-policy samples and off-policy samples from the replay buffer. We empirically determined an optimal 4:4 ratio for this through hyperparameter tuning, detailed in Supplementary Figure 7. To correct for policy drift in off-policy data, we apply importance sampling. The full objective for AVATAR combines the standard on-policy GRPO loss (Equation 2) with a corrected off-policy term:
$\mathcal{J}_{\text{AVATAR}}(\theta)=\mathcal{J}_{\text{on-policy}}(\theta)+\alpha\cdot\mathcal{J}_{\text{off-policy}}(\theta),$ (3)
where $\alpha$ controls the off-policy contribution. The off-policy objective $\mathcal{J}{\text{off-policy}}$ follows the same form as Equation 2, but uses an importance sampling ratio based on the behavior policy $\pi{\theta_{\text{off}}}$ that generated the replayed samples:
@@
Prior works [repo, vlrethinker] use generic replay strategies or standard advantage calculations for off-policy data whereas AVATAR introduces a difficulty-aware stratified buffer for curriculum learning and employs the VCRS baseline (Eq. 11) to stabilize advantage estimation specifically for these replayed samples. The full procedure is detailed in Algorithm 1.
4.2 Temporal Advantage Shaping (TAS)
Standard GRPO employs uniform credit assignment, applying the same sequence-level advantage $A_{i}$ to every token (Eq. 2), which struggles to differentiate crucial reasoning steps in complex tasks. Recent work [parthasarathi2025grpolambda, xie2025capo] demonstrates that not all tokens contribute equally: GRPO-$\lambda$ [parthasarathi2025grpolambda] shows that rapid value propagation to earlier tokens significantly improves learning, while process-based credit assignment [xie2025capo] reveals that tokens at critical decision points warrant stronger learning signals. For audio-visual reasoning requiring tight cross-modal coupling, this token-level differentiation becomes even more critical, like humans who devote disproportionate cognitive effort to planning how to align modalities and synthesizing multimodal evidence, rather than processing intermediate unimodal features. Motivated by these findings, we introduce Temporal Advantage Shaping (TAS), a simple, critic-free modification to GRPO. Reasoning often follows a planning-evidence-synthesis structure, reinforced during SFT. TAS leverages this by hypothesizing that planning and synthesis tokens are most critical. It modulates the learning signal based on relative token position, upweighting the sequence’s beginning and end, aligning with findings favoring non-uniform weighting [parthasarathi2025grpolambda]. As shown in Figure 2, TAS uses a U-shaped parabolic function $w_{t}$ to yield a token-specific shaped advantage, $A_{i,t}^{TAS}=w_{i,t}\cdot A_{i}$. This focuses learning on key reasoning phases, reinforcing critical steps without overemphasizing intermediate tokens.
Figure 2: To address GRPO’s uniform credit assignment (gray line), TAS applies a parabolic weighting function to amplify advantages during crucial planning and synthesis stages.
Parabolic Weighting Function.